# Flexible Tails for Normalizing Flows

### v1

@@ 4 Experiments
This section contains our empirical experiments. Firstly, recall that one motivation for our work in Section 1 is the claim that neural network optimisation can perform poorly under heavy tailed inputs. We verify this empirically in Section 4.1 for a simple neural network regression example, and then in Section 4.2 for normalizing flows. Secondly, Sections 4.24.4 contain normalizing flow examples, comparing our method to existing approaches for density estimation and a proof-of-concept variational inference example. Additional implementation details are provided in Appendix D.
4.1 Neural Network Regression with Extreme Inputs
This experiment considers the following simple regression problem
$\{X_{i}\}_{i=1}^{d}\sim t_{\nu},\quad Y\sim\mathcal{N}(X_{d},1).$
The $d$-dimensional input is heavy tailed. The output equals one of the inputs plus Gaussian noise.
Experimental Details
We consider a number of tail weight $(\nu$) and dimension ($d$) combinations, generating $5000$ train, validation and test samples for each. Our models are simple 2 layer multi-layer perceptrons with $50$ nodes in each hidden layer. The models are optimised with Adam to minimise mean square error, selecting the model with smallest validation loss found during training. We consider both sigmoid and ReLU activations. This is to investigate whether the main issue is saturation of sigmoid activation functions. The experiment was repeated 5 times, each trial sampling a new set of data.
Results
$d$ $\nu$ Sigmoid activation ReLU activation
30.0 1.01 (1.02) 1.01 (1.04)
5 2.0 1.74 (6.19) 1.06 (1.08)
1.0 1.57e+03 (8.57e+06) 3.45 (1.93e+03)
30.0 1.02 (1.02) 1.02 (1.05)
10 2.0 2.04 (7.28) 1.03 (1.05)
1.0 4.13e+03 (1.09e+06) 8.41 (49.6)
30.0 1.01 (1.04) 1.08 (1.09)
50 2.0 3.09 (5.26) 1.27 (12.3)
1.0 1.03e+04 (1.73e+06) 55.8 (6.59e+03)
30.0 1.02 (1.07) 1.15 (1.2)
100 2.0 3.41 (4.33) 1.60 (1.74)
1.0 1.02e+05 (1.94e+05) 336 (1.2e+03)
Table 1: Results of the neural network regression example. Values are the median test MSE over 5 trials, with the maximum provided in brackets. (We use median as it’s more robust than mean to extreme outliers. Maximum illustrates the presence and magnitude of such outliers.)
@@ 4.2 Density Estimation with Synthetic Data
In this model the only non-trivial dependency to learn is between $X_{d}$ and $X_{d-1}$, but there are also several heavy tailed nuisance variables.
We use this example to investigate whether modelling tails in the final layer of a NF is superior to modelling the tails in the base distribution. We look at how performance varies with changes in dimensionality $d$ and tail weight $\nu$. Section 4.1 showed that neural network regression can perform poorly under heavy tailed inputs. Here we investigate whether a similar finding holds in the setting of normalizing flows.
Some of the NF methods we compare involve fixed tail weight parameters (either $\nu$ for a Student’s T base distribution or $\lambda_{+},\lambda_{-}$ for our transformation), and for this experiment we fix these to their known true values. This means our analysis is not confounded by the difficulty of estimating these parameters. The tail weights are known as the marginal density for each $X_{i}$ is asymptotically proportional to $x_{i}^{-\nu-1}$. (This can be shown for $X_{d}$ by representing its density as a convolution and looking at the asymptotic behaviour.) Note that only $\lambda_{+},\lambda_{-}$ in (3) are fixed, not $\mu$ and $\sigma$.
@@
A baseline method, normal, uses a $d$-dimensional isotropic Gaussian base distribution. This is followed by an autoregressive rational quadratic spline layer then an autoregressive affine layer. The affine layer should be capable of capturing linear dependency, whilst the spline layer can adjust the shape of the body, but not the tails of the distribution.
Our proposed approach, tail transform flow, modifies the architecture just described by adding an additional layer for dealing with the tails, as described in Section 3. TTF trains the tail parameters alongside the other parameters. TTF (fix) is a 2-stage approach, which fixes the tail parameters to the known true values.
@@
As further variations on normal, we consider two alternative base distributions suggested by Amiri et al. (2022). The first is a Gaussian mixture model base distribution (m_normal) with 5 components (preliminary work found 10 or 15 components had no significant difference in performance). For each component we optimise the $d$ dimensional mean and diagonal covariance terms along with other parameters. The second base distribution is $d$ independent generalised normal distributions (g_normal). The generalised normal family does not have Pareto tails, so we cannot fix them to have the true tail weight. Instead, we optimise the shape parameter of each marginal.
We also consider COMET flows (COMET), as detailed in Section 1.1. For the normalizing flow part of this method we use the same architecture as normal. Our analysis is based on the code of McDonald et al. (2022). We don’t report results for $d=50$ as we found the code extremely slow to run in this case (discussed further in Appendix F). COMET is a 2-stage approach, which involves estimating tail parameters in the first stage, so again we fix these to the known correct values.
@@
We run 10 repeats for each flow/target combination. Each repeat samples a new set of data, with $5000$ observations. which is split in proportion 40/20/40, to give training, validation and test sets respectively. We train using the Adam optimiser with a learning rate of 5e-3. We use an early stopping procedure, stopping once there has been no improvement in validation loss in 100 epochs, and returning the model from the epoch with best validation loss. The loss plots from the optimisation were also visually inspected to confirm convergence. The selected model is then evaluated on the test set to give a negative test log likelihood per dimension, reported below for a range of dimensions and tail weights. Dividing by dimension acts to normalises the metric, allowing for easier comparison across dimensions.
$d$ Flow $\nu=0.5$ $\nu=1$ $\nu=2$ $\nu=30$
5 normal - - 2.01 (0.07) 1.46 (0.00)
m_normal - 403.24 (239.15) 1.94 (0.02) 1.46 (0.00)
g_normal - - 1.93 (0.01) 1.46 (0.00)
gTAF 6.42 (0.27) 2.49 (0.02) 1.90 (0.01) 1.46 (0.00)
TTF 3.33 (0.01) 2.34 (0.01) 1.89 (0.01) 1.47 (0.00)
mTAF 4.08 (0.03) 2.49 (0.02) 1.92 (0.01) 1.46 (0.00)
TTF (fix) 3.33 (0.01) 2.35 (0.01) 1.89 (0.01) 1.47 (0.00)
COMET 3.42 (0.01) 2.35 (0.01) 1.89 (0.00) 1.46 (0.00)
10 normal - - 2.00 (0.02) 1.46 (0.00)
m_normal - - 2.04 (0.07) 1.46 (0.00)
g_normal - - 1.98 (0.02) 1.46 (0.00)
gTAF 7.13 (0.31) 2.63 (0.02) 1.95 (0.00) 1.47 (0.00)
TTF 3.55 (0.01) 2.47 (0.00) 1.93 (0.00) 1.47 (0.00)
mTAF 4.48 (0.04) 2.63 (0.01) 1.95 (0.00) 1.46 (0.00)
TTF (fix) 3.54 (0.01) 2.46 (0.00) 1.93 (0.00) 1.47 (0.00)
COMET 3.63 (0.01) 2.46 (0.00) 1.93 (0.00) 1.47 (0.00)
50 normal - - 2.02 (0.01) 1.47 (0.00)
m_normal - - 2.02 (0.00) 1.47 (0.00)
g_normal - - 2.01 (0.00) 1.47 (0.00)
gTAF 7.49 (0.38) 2.65 (0.01) 1.99 (0.00) 1.47 (0.00)
TTF 3.68 (0.00) 2.54 (0.00) 1.98 (0.00) 1.47 (0.00)
mTAF 5.22 (0.04) 2.62 (0.01) 1.98 (0.00) 1.47 (0.00)
TTF (fix) 3.68 (0.00) 2.54 (0.00) 1.98 (0.00) 1.47 (0.00)
Table 2: Density estimation results on synthetic example. Each entry is a mean value of negative test log likelihood per dimension across 10 repeated experiments, with the standard error in brackets. Bold indicates methods whose mean log likelihood differs from the best mean by less than 2 standard errors (of the best mean). A dash indicates that the method did not converge (output was Inf or NaN) in at least one repeat.
@@ Results
Table 2 shows the results. For near-Gaussian tails ($\nu=30$) all methods perform similarly. The methods which are not specifically designed to permit GPD tails (normal, m_normal, g_normal) are the worst performing for $\nu\leq 2$, often not converging at all.
Fewer methods produce competitive results for heavier tails (small $\nu$) and higher dimensions (large $d$). Only TTF, TTF (fix) and mTAF are always competitive. COMET is also competitive for $d<50$, however it was not feasible to run it for $d=50$. In general COMET outperforms mTAF, and the TTF methods outperform COMET. We did not detect a significant difference between the two TTF methods.
Overall, the results suggest that it is advantageous to model tails in the final transformation (TTF, TTF (fix), COMET). Another question is whether it’s better to fix the tail parameters or optimise them. We found no significant difference here for TTF methods. However, since mTAF outperforms gTAF, fixing tail parameters is better for heavy tailed base distribution methods.
4.3 Density Estimation with S&P 500 Data
This section considers financial return data as an example of high dimensional multivariate data with extreme values. Our experiments target a varying number of dimensions $d\in\{100,200,300\}$, with data covering the time period 2010-01-04 to 2022-10-27, corresponding to 3227 days in total. For a given $d$, we take the closing prices of the top $d$ most traded S&P 500 stocks, and convert them in standard fashion to log returns i.e. the log returns are $x_{j}=\log(\frac{s_{j+1}}{s_{j}})$ where $s_{j}$ is stock closing price on day $j$. In this example we concentrate on the tails of the data, rather than time series structure. As such, we treat each day of log returns as an independent observation in $\mathbb{R}^{d}$. We normalize the log returns to zero mean and unit variance using the estimated mean and variance from the training and validation sets, as is standard practice in NF density estimation. For example, see the experimental framework of Papamakarios et al. (2017) and Durkan et al. (2019).
Preliminary analysis confirmed that the resulting data does exhibit heavy tails in the marginal distributions.
@@ 4.3 Density Estimation with S&P 500 Data
In this experiment we compare some NF architectures which performed reasonably well in Section 4.2: TTF, TTF (fix), mTAF and gTAF. It wasn’t feasible for us to run COMET flows for a problem of this dimension.
We reuse the NF architectures described in Section 4.2 with a slight alteration. We add trainable linear layers based on the LU factorisation (Oliva et al., 2018). For TTF methods, this immediately precedes the TTF layer. Otherwise, this is the final layer. The LU layers were included as we found that they greatly improved empirical performance.
See Section 3.4 for details of our initial stage of tail parameter estimation. Equivalent tail parameters are used for mTAF and TTF (fix). In some cases, this estimation procedure finds that a marginal distribution is light tailed. In these cases, as mentioned in Section 3.2, we set the tail parameter to a small value, corresponding to a degree of freedom of $\nu=1000$.
@@
We run 10 repeats for each flow/target combination. Each trains for 400 epochs using the Adam optimiser with a learning rate of 5e-4. The test set is comprised of observations after 2017-09-14, with train and validation sampled uniformly from the period up to and including this date. This corresponds to 1292 training, 645 validation and 1290 test observations respectively. We perform a standard early stopping procedure, with the final model chosen as that which achieved the lowest loss on the validation set.
@@ Results
The results are presented in Figure 2. As in Section 4.2, the TTF methods outperform the heavy tailed base distribution methods, although the advantage of TTF over mTAF is small. However, now fixing the tail parameters is advantageous for all methods: TTF (fix) significantly outperforms TTF, and mTAF outperforms gTAF.
Figure 2: Box plot of test negative log likelihoods per dimension from fitting S&P 500 log returns (10 repeats). Results above the dashed line use a two-stage procedure.
@@ 4.4 Variational Inference for Artificial Target
Here we consider variational inference using the target distribution (4). We previously used this in Section 4.2 to generate synthetic data for density estimation. As in that experiment, this model allows us to easily vary both target tail weights ($\nu$) and number of nuisance variables ($d$).
@@
We compare four flows which exhibited good performance in Section 4.2: TTF, TTF fix, mTAF and gTAF. COMET flows also performed well in the density estimation setting. However they involve making kernel density estimates of marginal distributions, which has no obvious equivalent for the VI setting. So we do not include them in this comparison. As in Section 4.2, for TTF fix and mTAF we fix the tail parameters to the known correct values.
@@
We consider $\nu\in\{1,2,30\}$, which covers; Cauchy, lighter than Cauchy and close to Gaussian nuisance noise respectively. (Unlike Section 4.2, we do not consider $\nu=0.5$ as $\nu=1$ already produces a clear conclusion of a single best method.) Additionally, we consider $d\in\{5,10,50\}$. We run 5 repeats for each flow/target combination. Each repeat uses 10,000 iterations of the Adam optimiser using a learning rate of 1e-3. The loss plots from the optimisation were visually inspected to confirm convergence. We use 100 Monte Carlo samples to produce gradient estimates. More details of the experiments are provided in Appendix D.
@@
We measure the accuracy of our fitted variational density using two diagnostics based on importance sampling and described in Appendix E: $ESS_{e}$ and $\hat{k}$.
@@ Results
Table 3 reports our results. All methods aside from gTAF achieve reasonable performance in terms of $\hat{k}$ (below $0.7$) in all settings. In terms of $ESS_{e}$, for $\nu=30$ the best results are for mTAF, but all methods do well ($ESS_{e}>0.7$). For heavier tails ($\nu\leq 2$), TTF fix has the best results in every case (with $ESS_{e}$ at least $0.75$). TTF produces slightly worse $ESS_{e}$ values (with $ESS_{e}$ at least $0.5$). The values for mTAF and gTAF are worse than either TTF method, especially for the $\nu=1,d=50$ case ($ESS_{e}<0.1$).
Overall the results show that, as before, it is advantageous to model tails in the final transformation (TTF, TTF fix). Fixing the tail parameters improves performance of TTF fix over TTF, but this is not consistently the case when comparing mTAF and gTAF.
$d$ Flow $\nu=1$ $\nu=2$ $\nu=30$
$ESS_{e}$ $\hat{k}$ $ESS_{e}$ $\hat{k}$ $ESS_{e}$ $\hat{k}$
5 gTAF 0.53 0.93 0.90 0.79 0.97 0.53
TTF 0.83 0.40 0.89 0.43 0.98 0.18
mTAF 0.10 0.18 0.52 -0.18 0.99 0.31
TTF (fix) 0.97 0.37 0.98 0.35 0.98 0.19
10 gTAF 0.28 1.05 0.74 0.89 0.96 0.28
TTF 0.90 0.36 0.92 0.36 0.96 0.20
mTAF 0.09 0.31 0.52 -0.37 0.98 0.25
TTF (fix) 0.93 0.33 0.95 0.32 0.97 0.07
50 gTAF 0.08 0.82 0.39 0.83 0.84 0.13
TTF 0.57 0.25 0.61 0.24 0.79 0.14
mTAF 0.02 0.57 0.41 0.03 0.90 0.17
TTF (fix) 0.75 0.18 0.81 0.18 0.85 0.08

### v2

@@ 4 Experiments
This section contains our experiments. Firstly, recall that one motivation for our work in Section 1 is the claim that neural network optimisation can perform poorly under heavy tailed inputs. We verify this empirically in Section 4.1 for NFs. This is also verified in Appendix F for a simple neural network regression example. Secondly, Sections 4.14.3 contain normalizing flow examples, comparing our method to existing approaches for density estimation and a proof-of-concept variational inference example. Additional implementation details are provided in Appendix G.
@@ 4.1 Density Estimation with Synthetic Data
In this model the only non-trivial dependency to learn is between $X_{d-1}$ and $X_{d}$, but there are also several heavy tailed nuisance variables.
We use this example to investigate whether modelling tails in the final layer of a NF is superior to modelling the tails in the base distribution. Of particular interest is how performance varies with dimensionality $d$ and tail weight $\nu$. Appendix F shows that neural network regression can perform poorly under heavy tailed inputs. Here we investigate whether a similar finding holds in the setting of NFs.
Some NF methods we test involve fixed tail weight parameters (either $\nu$ for a Student’s T base distribution or $\lambda_{+},\lambda_{-}$ for our transformation), and for this experiment we fix these to their known true values. This means our analysis is not confounded by the difficulty of estimating these parameters. The tail weights are known since the marginal density for each $X_{i}$ can be shown to be asymptotically proportional to $x_{i}^{-\nu-1}$. Note that only $\lambda_{+},\lambda_{-}$ in (3) are fixed, not $\mu$ and $\sigma$.
@@
A baseline method, normal, uses a $d$-dimensional isotropic Gaussian base distribution. This is followed by an autoregressive rational quadratic spline (RQS) layer then an autoregressive affine layer. The latter should be capable of capturing linear dependency, while the former can adjust the shape of the body, but not the tails of the distribution.
Our proposed approach, tail transform flow, modifies the architecture just described by adding an additional layer for dealing with the tails, as described in Section 3. TTF trains the tail parameters alongside the other parameters. TTFfix is a 2-stage approach, which fixes the tail parameters to the known true values.
@@
As further variations on normal, we consider two alternative base distributions suggested by Amiri et al. (2022) – Gaussian mixture (m_normal) and Generalised Normal (g_normal). More details are provided in Appendix G.
We also consider COMET flows (COMET), as detailed in Section 1.1. For the normalizing flow part of this method we use the same architecture as normal. Our analysis is based on the code of McDonald et al. (2022) with some improvements to implementation details (needed to run more complicated examples later). COMET is a 2-stage approach, which involves estimating tail parameters in the first stage, so again we fix these to the known correct values.
We also tested a variant suggested by a reviewer, TTF_tBase. This combines TTF with a Student’s $T$ base distribution with trainable degrees of freedom. It performed worse than both TTF and TTFfix, so is omitted from the main paper results for brevity. Its results appear in Appendix J.1.
@@
We run 10 repeats for each flow/target combination. Each repeat samples a new set of data, with $5000$ observations. which is split in proportion 40/20/40, to give training, validation and test sets respectively. We train using the Adam optimiser with a learning rate of 5e-3. We use an early stopping procedure, stopping once there has been no improvement in validation loss in 100 epochs, and returning the model from the epoch with best validation loss. Optimisation loss plots were also visually inspected to confirm convergence. The selected model was then evaluated on the test set to give a negative test log likelihood per dimension.
Table 1: Density estimation results on synthetic example for $d=50$. Each entry is a mean value of negative test log likelihood per dimension across 10 repeated experiments, with the standard error in brackets. Bold indicates methods whose mean log likelihood differs from the best mean by less than 2 standard errors (of the best mean). A dash indicates potential unstable optimisation (at least one repeat had a final loss above 1e5).
Flow $\nu=0.5$ $\nu=1$ $\nu=2$
normal - - 2.02 (0.01)
m_normal - - 2.02 (0.00)
g_normal - - 2.01 (0.00)
gTAF 7.49 (0.38) 2.65 (0.01) 1.99 (0.00)
TTF 3.68 (0.00) 2.54 (0.00) 1.98 (0.00)
mTAF 5.22 (0.04) 2.62 (0.01) 1.98 (0.00)
TTFfix 3.68 (0.00) 2.54 (0.00) 1.98 (0.00)
COMET 3.74 (0.00) 2.55 (0.00) 1.97 (0.00)
@@ Results
Table 1 shows a selection of results with $d=50$. See Appendix J.1 for other $d$ and $\nu$ values. The methods not specifically designed to permit GPD tails (normal, m_normal, g_normal) are the worst performing, often not converging at all for $\nu\leq 1$. However, the difference between methods is small for $\nu=2$, and all methods are similar for $\nu=30$ (near-Gaussian tails—results in appendix).
For particularly heavy tails $(\nu\leq 1)$, the best performing methods are TTF, TTFfix and COMET: the approaches which model tails in the final transformation. We did not detect a significant difference between the two TTF methods, but they both outperform COMET. A similar pattern is present for other choices of $d$ (results in appendix).
It is interesting that COMET performs competitively here, despite not inducing Fréchet tails. A possible reason is that COMET can produce log-normal tails (see Appendix C.2), and these have similar sub-asymptotic properties to GPD tails (Nair et al., 2022).
Another question is whether it’s better to fix the tail parameters or optimise them. Table 1 shows that mTAF outperforms gTAF, so fixing tail parameters is better for heavy tailed base distribution methods. However there is no significant difference between TTF methods. A further investigation of the results suggested by a reviewer does suggest possible advantages of TTF which might become more significant in other settings. See Appendix J.2 for details.
4.2 Density Estimation with Real Data
This section investigates density estimation for several real datasets with extreme values, covering insurance, financial and weather applications. Three are taken from Liang et al. (2022); Laszkiewicz et al. (2022) and one is novel to this paper. Appendix I has more information about the datasets and standard preprocessing applied before density estimation.
@@ 4.2 Density Estimation with Real Data
In this experiment we compare NF architectures which performed reasonably well in Section 4.1: TTF, TTFfix, mTAF, gTAF and COMET.
For most examples we reuse the NF architectures described in Section 4.1 with a slight alteration. We add trainable linear layers based on the LU factorisation (Oliva et al., 2018). For TTF methods, this immediately precedes the TTF layer. Otherwise, this is the final layer. We found adding LU layers greatly improved empirical performance.
For the most complex dataset (CLIMDEX) we use a deeper architecture, as used in Laszkiewicz et al. (2022), to permit fair comparison with their results. This has 5 RQS layers, alternated with LU layers. We use this architecture for mTAF and gTAF, and modify it for TTF, TTFfix and COMET in the same way as described in Section 4.1.
@@
For all examples we run 10 repeats for each flow/target combination. For most examples we train for 400 epochs using the Adam optimiser with a learning rate of 5e-4. The exception is CLIMDEX, where we follow the more complex training setup of Laszkiewicz et al. (2022) (e.g. cosine annealing, more optimisation steps) to permit fair comparison with their results.
@@ Results
The results are presented in Table 2. TTF outperforms previous NF methods for density estimation: it is the best performing method for 3 examples, and second best to TTFfix for the other. The performance of TTFfix is more variable: it is the best performing in one example, second best to TTF in two examples, but the worst performing in the remaining example.
Overall these results show TTF methods provide an improvement in density estimation performance, and that the additional freedom of TTF to learn tail behaviour compared to TTFfix (e.g. it can learn tail asymmetry) is beneficial in some cases, especially the most complicated dataset, CLIMDEX.
Also, estimating tail parameters for each marginal involves extra compute costs, especially for high dimensional datasets such as CLIMDEX, where the cost was comparable to fitting the normalizing flow (although easy to parallelise).
Table 2: Real data density estimation results. Each entry is a mean test negative log likelihood over 10 trials, with standard deviation reported in brackets. Values marked with * are those reported by Laszkiewicz et al. (2022). Results within one standard deviation of the best mean are highlighted in bold.
Model Insurance Fama 5 S&P 500 CLIMDEX
normal 1.41 (0.03) 4.65 (0.01) 334.01 (1.02) -2101.91 (9.44)*
gTAF 1.41 (0.03) 4.68 (0.01) 321.81 (0.55) -2113.48 (7.93)*
TTF 1.37 (0.02) 4.61 (0.01) 317.56 (0.56) -2214.28 (13.25)
mTAF 1.52 (0.03) 4.90 (0.03) 322.98 (0.33) -2121.38 (10.91)*
TTFfix 1.38 (0.01) 4.63 (0.01) 314.84 (0.46) -2090.91 (11.90)
COMET 1.41 (0.03) 4.63 (0.01) 324.38 (0.58) -2118.60 (8.61)
@@ 4.3 Variational Inference for Artificial Target
As a proof-of-concept investigation of variational inference, we perform experiments using (4) as a target distribution. Recall that in Section 4.1 we used this to generate synthetic data for density estimation. This model allows us to easily vary tail weights ($\nu$) and number of nuisance variables ($d$).
@@
We compare four flows which exhibited good performance in Section 4.1: TTF, TTFfix, mTAF and gTAF. While COMET flows also performed well in the density estimation setting, they involve making kernel density estimates of marginal distributions, which has no obvious equivalent for the VI setting. So we do not include them in this comparison. As in Section 4.1, for TTFfix and mTAF we fix the tail parameters to the known correct values.
@@
We take $d\in\{5,10,50\}$ and $\nu\in\{0.5,1,2,30\}$ (heavier than Cauchy; Cauchy; lighter than Cauchy; close to Gaussian). We run 5 repeats for each flow/target combination. Each repeat uses 10,000 iterations of the Adam optimiser with learning rate 1e-3. Optimisation loss plots were visually inspected to confirm convergence.
@@
We measure the accuracy of our fitted variational density using two diagnostics based on importance sampling and described in Appendix H: $ESS_{e}$ and $\hat{k}$.
@@ Results
Table 3 reports our results. TTFfix has best $ESS_{e}$ for heavier tails ($\nu\leq 2$). TTF produces slightly worse values, with mTAF and gTAF worse than either TTF method, especially for the $\nu\leq 1,d=50$ cases ($ESS_{e}<0.1$). For $\nu=30$ the best results are for mTAF, but all methods do well ($ESS_{e}>0.7$). All methods aside from gTAF achieve useful approximations ($\hat{k}$ below $0.7$), in every setting where $\nu\geq 1$. In the most challenging setting—$\nu=0.5,d=50$—both TTF method achieve good $\hat{k}$ values.
Overall the results show that, as before, it is advantageous to model tails in the final transformation (TTF, TTFfix). Fixing tail parameters improves performance of TTFfix over TTF, perhaps because optimisation is more stable in this case.
@@
$d$ Flow $\nu=0.5$ $\nu=1$ $\nu=2$ $\nu=30$
$ESS_{e}$ $\hat{k}$ $ESS_{e}$ $\hat{k}$ $ESS_{e}$ $\hat{k}$ $ESS_{e}$ $\hat{k}$
5 gTAF 0.47 0.89 0.53 0.93 0.90 0.79 0.97 0.53
TTF 0.59 0.67 0.83 0.40 0.89 0.43 0.98 0.18
mTAF 0.00 2.25 0.10 0.18 0.52 -0.18 0.99 0.31
TTFfix 0.62 0.82 0.97 0.37 0.98 0.35 0.98 0.19
10 gTAF 0.19 0.64 0.28 1.05 0.74 0.89 0.96 0.28
TTF 0.40 0.74 0.90 0.36 0.92 0.36 0.96 0.20
mTAF 0.00 3.13 0.09 0.31 0.52 -0.37 0.98 0.25
TTFfix 0.45 0.88 0.93 0.33 0.95 0.32 0.97 0.07
50 gTAF 0.02 0.90 0.08 0.82 0.39 0.83 0.84 0.13
TTF 0.39 0.51 0.57 0.25 0.61 0.24 0.79 0.14
mTAF 0.00 7.92 0.02 0.57 0.41 0.03 0.90 0.17
TTFfix 0.43 0.53 0.75 0.18 0.81 0.18 0.85 0.08